Spartan Credit Bank is offering 7.6 percent compounded daily on its savings accounts. You deposit $6,000 today.


a. How much will you have in the account in 5 years? (Use 365 days a year. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
b.

How much will you have in the account in 10 years? (Use 365 days a year. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

c. How much will you have in the account in 20 years? (Use 365 days a year. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/365)^365n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.A=6000*(1+0.076/365)^(365*5)

=6000*1.462226748

=$8773.36(Approx).

2.A=6000*(1+0.076/365)^(365*10)

=6000*2.138107063

=$12828.64(Approx).

3.A=6000*(1+0.076/365)^(365*20)

=6000*4.571501814

=$27429.01(Approx).